a television and a radio together cost $720. The television is $400 dearer than the radio. how much does the television cost?
Answer by Cromlix(4381)

Television = x + 400
Radio = x
x + (x + 400) = 720
2x + 400 = 720
2x = 720 - 400
2x = 320
x = 160
Therefore television costs x + 400
160 + 400
$560
radio costs x
costs $160
Hope this helps
:-)